You need to make a password for you computer that contains two letters and two numbers that cannot repeat. How many different passwords are possible?
Answer by jrfrunner(365)

You need passwords with WWNN without repeating the letters or numbers
---
so there are 26 words to choose from, after the first word, there are 25 words left to choose. There are 10 one digit number 0 to 9, but after you select the first number there are 9 choices left.
--
So WWNN = 26*25*10*9 = 58500, but.... (and there is always a but)
--
there are 4! ways to rearrange these 2 letters adn 2 numbers, creating different passwords from the "same" letters and numbers.
--
so the number of different passwords is 4!*26*25*10*9=1404000